Observers: Tomás Müller, Juanita Antilen
Support Team: Phil Wiseman, Mat Smith, Matt Grayling, Chris Frohmaier
TAT: Janet Chen, Regis Cartier, Roberta Carini
EFOSC calibrations (biases, flats and arcs) done in the afternoon
Gr18+20 calibrations in the morning
Start: Humidity 39%, wind speed 2.9 m/s, seeing: 1.8".
Middle: Humidity 37%, wind speed 1.4 m/s, seeing: 0.8".
End : Humidity 34%, wind speed 2.3 m/s, seeing: 1.0".
Cloudy at the beginning of the night. The telescope's dome was opened at ~00:40 UT.
The weather conditions for the rest of the night were mostly good.